• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kent, Pete

    Log entry time 14:34:04 on June28,2004

    Entry number 127241

    keyword=Left Profile Scanner Update

    Left profile scanner.
    
    LV to the potentiometers was changed from 9.5V (readout on a multimeter,but reading 8.7V on the front panel of power supply) to reading 8.54V on the voltmeter and 7.7 on the front panel.  
    
    The potentiometer was also rewired by Pete to make sure all of the
     connections were solid.
    
    While scanner was moving, Pete measured the voltage output from the
    Y potentiometer.  Voltages behaved normally, and still only pedestal
    was readout from upstairs.  Obviously the readout chain was the 
    problem.  With no scope to diagnose, we changed what we could.  We
    moved the signal up one channel on the V2F (from 3 -> 2), and changed
    the jumper from the V2F to match.  We also changed the level translator channel up one.  The effect of this is to move
    Lprofy from scaler5_7 to scaler5_6.  Also, we moved the cable
    carrying this signal to the spectrometer DAQ up to match, so this 
    should be entire transparent to the spectrometer.
    
    This worked; good readout is now seen from lprofy.  The problem was probably with the V2F channel 3, because the non-working state appeared to correspond to roughly the V2F pedestal.
    
    
    PMT noise issues were not addressed during this access.